OpenStudy (anonymous):

can sum1 pls help me!? the value of x satisfying √(x+1) - √(x-1) =1 is a. 5/4 b. 1/2

OpenStudy (anonymous):

\[\sqrt{x+1} - \sqrt{x-1}=1\]

OpenStudy (anonymous):

ok for this problem, we need to isolate one of the radical terms. Your first step would be to add root x-1 to both sides of the equation

OpenStudy (anonymous):

first square the equation x+1 + x-1 +2sqrt(x^2-1)=1 2x+2sqrt(x^2-1)=1 2sqrt(x^2-1)=1-2x square again 4(x^2-1)=1+4x^2-4x -4=1-4x 4x=5 x=5/4.

OpenStudy (anonymous):

tnx a lot...shud v do (a +b ) ^2 ?

OpenStudy (anonymous):

yeah (a+b)^2=a^2+b^2+2ab

OpenStudy (anonymous):

so will it be 2x +2\[\sqrt{x ^{2 -1}} = 1\]

OpenStudy (anonymous):

Or just directly substitute the options in the equation and see which satisfies... like if you substitue 1/2 you have a sqrt(x-1)=sqrt(-1/2) which is not defined so 1/2 is wrong now sqrt(5/4 -1) -sqrt(5/4 +1)= sqrt(9/4)-sqrt(1/4)=3/2 -1/2 =1 so 5/4 is the answer

OpenStudy (anonymous):

yep...thank u .. :)